OpenStudy (anonymous):

What is the general term formula of the following number series? 1, √2/2, 1/2, √2/4, 1/4.

OpenStudy (xapproachesinfinity):

First what is the first term and the ratio how did we get from 1 to \(\large \frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

Maybe 1*√2/2?

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

there is no may be here my friend. there is only it is or not hehe so to go from 1 to root2/2 you need to multiply by 1/root2 and then again from root2/2 you multiply by 1/root2 you 1/2 and so on therefore we call that number we are multiplying by the ratio of a geometric sequence the first term is 1 since we started with that number!

OpenStudy (xapproachesinfinity):

By now since you are taking this course you should be familiar with the general form of this sequence

OpenStudy (anonymous):

So, 1*√2/2=√2/2 √2/2*√2/2=2/4=1/2 1/2* √2/2=√2/4 √2/4*√2/2=2/8=1/4.

OpenStudy (anonymous):

Am I right?

OpenStudy (xapproachesinfinity):

Correct! that number root2/2 is the ratio

OpenStudy (xapproachesinfinity):

denoted r for geometric sequences

OpenStudy (anonymous):

And the general term formula is \[\left( √2/2\right)^{n-1}\]

OpenStudy (xapproachesinfinity):

Correct!

OpenStudy (anonymous):

Thank you!

OpenStudy (xapproachesinfinity):

just don't forget the first term in other examples in this case is 1 so i doesn't affect the form

OpenStudy (xapproachesinfinity):

You are wlm!